OpenStudy (anonymous):

Find the equation of the plane through (7,9,5) parallel to the plane 3x + 4y + 7z = 0.

OpenStudy (amistre64):

your gradient is <3,4,7> which is also your normal

OpenStudy (anonymous):

amistre knows her vectors well. :)

OpenStudy (amistre64):

3(x-7)+4(y-9)+7(z-5) = 0

OpenStudy (amistre64):

lol... at least 2 of them ;)

OpenStudy (amistre64):

another form of it is: 3x +4y + 7z - (21+36+35) = 0

OpenStudy (anonymous):

amistre is a her? :O

OpenStudy (amistre64):

online im whatever they wanna picture me as lol

OpenStudy (anonymous):

yeah she is a she

OpenStudy (anonymous):

or i projected her as a she. a cute she

OpenStudy (anonymous):

LOL. is that the eq?

OpenStudy (amistre64):

that is the eq

OpenStudy (amistre64):

i look like meg ryan if you squint hard enough ;)

OpenStudy (anonymous):

aweesome. haha howd u get 21 and those other numbers

OpenStudy (amistre64):

first i killed a chicken, then spread its blood around the bedposts, did an incantation, and oijied the results ...maybe

OpenStudy (anonymous):

wow thats some talent!

OpenStudy (amistre64):

take your normal/gradient; and apply it to the point they gave you

OpenStudy (anonymous):

haha lol aight

OpenStudy (amistre64):

x-7 y-...... z - ......

OpenStudy (amistre64):

i just exapnded the first equation i gave you; they are equal

OpenStudy (anonymous):

sweet

OpenStudy (amistre64):

for reference; the normal vector is <a,b,c>; and point P(x,y,z) the plane equation is: a(x-Px)+b(y-Py)+c(z-Pz)=0

OpenStudy (anonymous):

ohh awesome.got it!